Calculate the value of C3+C7+C11......

Grade:9

1 Answers

Arun Kumar IIT Delhi
10 years ago
Hello Student,

We know that
\\(1+x)^n \\=>C_1+C_3+C_5....=2^{n-1} \\$on putting x=i and equating real and imaginary parts$ \\C_1-C_3+C_5-C_7....=2^{n-1}sin{\n\pi \over 4} \\$subtracting we get $ \\ C_3+C_7+C_{11}....={2^{n-1}-2^{n-1}sin{\n\pi \over 4} \over 2}
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty